OpenStudy (fibonaccichick666):

Proof question: Can you please describe the process for proving the existence of a limit at infinity by using the definition of a limit?

OpenStudy (anonymous):

you mean \[\lim_{x\to \infty}f(x)=L\]?

OpenStudy (fibonaccichick666):

The definition is thus: Let \(f:S\rightarrow R\) be a function and c a cluster point of S. Suppose that there exists an \(L \in R\) and for every \(\epsilon >0, \)there exists a \(\delta >0\) such that whenever \(x\in S /{c} \)and \(|x-c|<\delta\), then \[\lim_{x\rightarrow c} f(x):=L\] OR \[f(x)\rightarrow L ~as ~x\rightarrow c\]. If no such limit exists we say the limit DNE.

OpenStudy (fibonaccichick666):

no, ok so my problem is \[\lim_{x\rightarrow c} \sqrt{x} ~~for~~c\ge 0\] So, I know the limit equals \(+\infty\) but I don't know how to show that

OpenStudy (anonymous):

because \(\sqrt{x}\) is continuous, so the limit is \(\sqrt c\)

OpenStudy (fibonaccichick666):

but really, I just need to work on proving by definition for limits. And that's not proof enough

OpenStudy (anonymous):

it is certainly not \(\infty\)

OpenStudy (anonymous):

it is not a proof, true, but at least it is right!

OpenStudy (fibonaccichick666):

c is greater than or equal to zero though

OpenStudy (fibonaccichick666):

so I took it one step farther

OpenStudy (fibonaccichick666):

you don't think it wants c to be a variable?

OpenStudy (anonymous):

there is no infinity involved here

OpenStudy (fibonaccichick666):

How come? If c is greater than zero and a variable wouldn't that imply a never ending set of values?

OpenStudy (anonymous):

it is asking you to prove \[\lim_{x\to 4}\sqrt x=2\] only with \(c\) instead of \(4\)

OpenStudy (fibonaccichick666):

ohhhh

OpenStudy (anonymous):

oooh i see what confused you!!

OpenStudy (anonymous):

the \(c>0\) is just saying what the domain is !!!

OpenStudy (fibonaccichick666):

yea, that just didn't process

OpenStudy (fibonaccichick666):

thanks sat

OpenStudy (anonymous):

yw glad to help, especially since it was so easy

OpenStudy (fibonaccichick666):

follow up, for proofs by definition, how does it work

OpenStudy (fibonaccichick666):

lol yea, i know brain farts eh?

OpenStudy (anonymous):

just do it as they say

OpenStudy (fibonaccichick666):

that's the problem... they don't say. (it's basic analysis by jiri liebel, essentially his lecture notes) I do not know the structure for this style of proof yet

OpenStudy (anonymous):

well that is not really correct, you have to work backwards your goal it to find the \(\delta\) that works for ever \(\epsilon\) in other words, \(\delta \) will be something like \(\frac{\epsilon}{2}\) or \(\epsilon^2\) or sommat

OpenStudy (anonymous):

wanna try it with a number?

OpenStudy (fibonaccichick666):

the example is using delta as a minimum of some weird function they pulled out of thin air

OpenStudy (fibonaccichick666):

yea

OpenStudy (fibonaccichick666):

if we could

OpenStudy (anonymous):

yeah some gimmick probably lets prove \[\lim_{x\to 4}\sqrt x=2\] as one of my favorite math teacher one said, "just because it is obvious, doesn't mean you can't prove it"

OpenStudy (fibonaccichick666):

haha tell me about it

OpenStudy (fibonaccichick666):

ok so, first step is to define our f(x) right?

OpenStudy (anonymous):

the first step is translating the general in to the specific now that i read carefully what you wrote these is something missing

OpenStudy (fibonaccichick666):

hm?

OpenStudy (anonymous):

somewhere you should have \[|f(x)-L|<\epsilon\]

OpenStudy (fibonaccichick666):

oh yea, I missed a line

OpenStudy (fibonaccichick666):

instead of the statement of what it looks like to be convergent that should be in it's place

OpenStudy (anonymous):

for every \(\epsilon\) there is a \(\delta\) such that if \(|x-c|<\delta\) then \(|f(x)-L|<\epsilon\)

OpenStudy (anonymous):

in our case \(f(x)=\sqrt{x}, c=4\)

OpenStudy (anonymous):

and \(L=2\)

OpenStudy (fibonaccichick666):

ok right so this is where I'm confused, how do we find delta?

OpenStudy (fibonaccichick666):

do we do |f(x)-√c|>delta?

OpenStudy (anonymous):

we work backwards as i said first we go from the general to the specific we want \[|\sqrt{x}-2|<\epsilon\]

OpenStudy (fibonaccichick666):

ok

OpenStudy (anonymous):

we can control \(|x-4|\) i.e. we can make it as small as we like, so in general what we want to look for is to somehow get an expression involving \(|x-4|\) out of the inequality \[|\sqrt{x}-2|<\epsilon\]

OpenStudy (fibonaccichick666):

so then we need to use the triangle ineq?

OpenStudy (anonymous):

nah lets do some simple algebra and see if something nice happens

OpenStudy (fibonaccichick666):

why |x-4|? How did you get that?

OpenStudy (anonymous):

because in our case \(c=4\)

OpenStudy (fibonaccichick666):

ok and since we know |x-c|<delta, that may be useful?

OpenStudy (anonymous):

recall the line \(|x-c|<\delta\)

OpenStudy (anonymous):

it isn't useful, it is essential we are trying to prove the limit as x goes to 4 is 2 without it, it is not true

OpenStudy (anonymous):

hold on a second lets see how this game is played you tell me that the limit as x goes to 4 of root x is 2

OpenStudy (fibonaccichick666):

ok, so will we always try to make the original have some form of |x-c|?

OpenStudy (fibonaccichick666):

k

OpenStudy (anonymous):

i say, ok really? make the square root of x to be within 0.1 of 2 you say, ok, i will make \(x=4.3\) because \(\sqrt{4.3}=2.07\) which is less that \(.1\) away from \(2\)

OpenStudy (fibonaccichick666):

ok i follow

OpenStudy (anonymous):

then i say, no no i didn't mean \(.1\) away from 2, that is too easy, how about \(.01\) away from 2 you reply no big deal lets make \(x=4.03\) because \(\sqrt{4.03}=2.007\) which is within \(.01\) like you asked for

OpenStudy (fibonaccichick666):

alright

OpenStudy (anonymous):

damn i say, how about making it \(.0001\) away from 2 and you come back with give me a break, i'll just make \(x=4.0003\) and get \(\sqrt{4.003}=2.0007\) so there!

OpenStudy (anonymous):

and so on

OpenStudy (fibonaccichick666):

ok, but I don't see how that correlates here

OpenStudy (anonymous):

so finally i just say ok wise guy , make \(\sqrt{x}\) within \(\epsilon\) units of \(2\) and i can make \(\epsilon\) as small as i like any number , \(.00001\) or \(10^{-40}\) or whatever

OpenStudy (anonymous):

and now you have to answer me by saying, ok i will make \(x\) this far from \(4\) where the "this far" is some expression in terms of \(\epsilon\)

OpenStudy (fibonaccichick666):

ok, I understand the question but am at a loss as to how to answer it

OpenStudy (anonymous):

in other words we leave the world of numberic examples and enter the higher realm of variables not \(.1\) or \(.001\) or \(10^{50}\) but \(\epsilon\) for ANY \(\epsilon\)

OpenStudy (anonymous):

well that is the tough part but usually it involves not more than some algebra, maybe a trick as well so lets see what we need to do

OpenStudy (fibonaccichick666):

okso we start with \(|\sqrt{x}-2|<\epsilon\)

OpenStudy (anonymous):

we need to force \(|\sqrt{x}-2|<\epsilon\) what does that really say?

OpenStudy (fibonaccichick666):

that says \(|\sqrt{x}|-|\sqrt{2}|<\epsilon\)

OpenStudy (fibonaccichick666):

b y triangle ineq

OpenStudy (anonymous):

it says \[2-\epsilon<\sqrt{x}<2+\epsilon\][

OpenStudy (fibonaccichick666):

that shouldn't be a square root on my two

OpenStudy (anonymous):

oh wait, i recall there is a gimmick for this one yeah i don't think we need the triangle

OpenStudy (anonymous):

but lets see what we get without the gimmick

OpenStudy (fibonaccichick666):

ok

OpenStudy (fibonaccichick666):

should we square everything?

OpenStudy (anonymous):

yeah lets

OpenStudy (anonymous):

\[4-4\epsilon+\epsilon^2<x<4+4\epsilon+\epsilon^2\] now some more algebra to find out \(|x-4|\)

OpenStudy (fibonaccichick666):

so we get \(4-4\epsilon+\epsilon ^2<x<4+4\epsilon+\epsilon^2\)

OpenStudy (fibonaccichick666):

\(x-4<4\epsilon +\epsilon^2\)

OpenStudy (fibonaccichick666):

?

OpenStudy (anonymous):

then \[-4\epsilon+\epsilon^2<x-4<4\epsilon+\epsilon^2\] if it wasn't for that damed \(\epsilon^2\) we would be done

OpenStudy (fibonaccichick666):

can we just factor it out?

OpenStudy (anonymous):

we would have \[-4\epsilon<x-4<4\epsilon\] or \[|x-4|<4\epsilon\]

OpenStudy (anonymous):

and we would make our \(\delta=4\epsilon\) and be finished

OpenStudy (fibonaccichick666):

wait, since we are working ineqs we would just need to show that 4\epsilon is greater than or equalt to 4\epsilon +\epsilon^2

OpenStudy (fibonaccichick666):

before we could drop it off... right?

OpenStudy (anonymous):

hmm maybe i have to think about that seems reasonable

OpenStudy (fibonaccichick666):

nvm, it's have to be = not ge

OpenStudy (fibonaccichick666):

can we use that epsilon goes to zero?

OpenStudy (anonymous):

in any case, that is how these proof work i think the case you have the usual gimmick in showing that \[\lim_{x\to c}\sqrt{ x}=\sqrt{c}\]is to take \[|\sqrt{x}-\sqrt{c}|\] and multiply top and bottom by the conjugate

OpenStudy (anonymous):

then you bound the denominator

OpenStudy (fibonaccichick666):

ah ok, and I would want |x-c| for my malleable number

OpenStudy (anonymous):

you get \[\frac{|\sqrt{x}-\sqrt{c}|}{|\sqrt{x}+\sqrt{c}|}\times |\sqrt{x}+\sqrt{c}|\] \[=\frac{|x-c|}{\sqrt{x}+\sqrt{c}}\]

OpenStudy (anonymous):

right exactly you have control on how small you can make \(|x-c|\)

OpenStudy (fibonaccichick666):

then move the |x-c| over to the epsilon side or no?

OpenStudy (anonymous):

you want \(\frac{|x-c|}{\sqrt{x}+\sqrt{c}}<\epsilon\) after doing the algebra,

OpenStudy (fibonaccichick666):

no, no, move the denominator over so that we get our delta value right?